• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Xinhu Yan

    Log entry time 15:58:57 on December17,2007

    Entry number 221666

    keyword=Day shift Summary

    Day shift summary
    SL:Xinhu Yan
    TO:Konrad Aniol
    3rd:Yunxiu Ye


    8:00 Day shift starting ,but we were waiting for Loop2 stabliztion,Pass changed and computer problem fixed
    10:20-11:10 Bob came here to reboot the adaqs and got the DAQ back to normal ,now we were taking the cosmic run and waiting for the pass changed
    11:44 Moved the spectrometer from 23 degree to 15 degree
    12:00 the beam energy was 2.13938Gev and changed the momentum to 2.066278Gev/c for left and right spectrometer
    12:30 beam had come back ,checked the OTR ,compton and etc.
    13:48 Harp scan for the beam
    16:00 Day shift end


    ---------------------------------------------------------------------------------------
    Left Spectrometer
    ----------------------------------------------
    E0=2.13938Gev 15 degree
    p0=2.066278Gev/c
    Run 4164 Cosmic run


    ----------------------------------------------------------------------------------------
    Right Spectrometer
    ----------------------------------------------
    E0=2.13938Gev 15 degree
    p0=2.066278Gev/c
    Run 22049 Cosmic run